OpenStudy (anonymous):

how to calculate direction of 4d vector

OpenStudy (anonymous):

Divide it by its magnitude.

OpenStudy (wolfe8):

^ which is the same as finding the direction of any vector in any dimension if I am right

OpenStudy (anonymous):

for example <1,2,3,4> what to do

OpenStudy (anonymous):

i m assuming \[\frac{ 1 }{ 30 }<1,2,3,4>\] will be direction of this vector??? @wio

OpenStudy (anonymous):

The magnitude is 30?

OpenStudy (anonymous):

oops sqrt(30)

OpenStudy (anonymous):

That looks right.

OpenStudy (anonymous):

\[ \frac{\sqrt{30}}{30} \]

OpenStudy (anonymous):

can yoou explain me graphicaly how is this a direction and related to vector?

OpenStudy (anonymous):

We have to use a unit vector to describe the direction of a vector, since it takes away the magnitude of the vector

ganeshie8 (ganeshie8):

you will need to use 4 directional cosines to describe the directio i think, like u do in 3d

OpenStudy (anonymous):

ohk...so basically unit vector is the direction of vector v

OpenStudy (anonymous):

oh how? so unit vector isn't right?

ganeshie8 (ganeshie8):

unit vector is the vector wid magnitude of 1, along the direction of given vector.

OpenStudy (anonymous):

so it is direction, isnt it!!! i m confused

ganeshie8 (ganeshie8):